Jefferson Lab HALL A Logbook for December 14 - All EXCEPT Automatic Entries

  • Main INDEX

    Sunday

    
    

    Saturday

    
    

    Friday

    
    

    Thursday

    
    

    Wednesday

    
    

    Tuesday

    
    

    Monday

    
    

    Sunday

    
    

    Saturday

    
    

    Friday

    
    

    Thursday

    
    

    Wednesday

  • 404743 12/10/14 15:43 vmgray Day Shift Target screen shot
  • Tuesday

  • 404726 12/09/14 15:55 Hall A (Tech Checklist
  • 404709 12/09/14 08:51 R. Michaels problems seen in online analysis
  • Monday

  • 404695 12/08/14 00:31 J. Roche BCM Calibration (unser versus bcms)
  • 404694 12/08/14 00:01 Hashir Target Screenshot
  • Sunday

  • 404676 12/07/14 18:40 A. Camsonne HARP scan test
  • 404675 12/07/14 18:29 Hashir Target Moved
  • 404674 12/07/14 18:03 Hashir Target moved
  • 404673 12/07/14 18:01 Hashir Target Moved
  • 404671 12/07/14 17:08 Hashir Target Screenshots
  • 404670 12/07/14 17:07 Hashir Target Screenshot
  • 404668 12/07/14 15:36 gautam Target screenshot
  • 404663 12/07/14 13:15 gautam Target screenshot
  • 404652 12/07/14 08:34 gautam target screen and charts
  • 404648 12/07/14 07:49 Tyler Hague Target and strip charts
  • 404645 12/07/14 04:26 Tyler Hague IOC reboot - JT valve communication failure
  • 404644 12/07/14 03:30 Tyler Hague JT valve control - IOC reboot
  • Saturday

  • 404642 12/06/14 22:38 N. Israel target and charts
  • 404612 12/06/14 12:21 wdconinc Target status screenshots
  • 404578 12/06/14 03:34 akabir Encoder positions
  • Friday

  • 404574 12/05/14 23:58 N. Israel target charts
  • 404572 12/05/14 20:48 N. Israel target screen and charts
  • 404561 12/05/14 15:55 blomadam Target computer beam current readout broken
  • 404560 12/05/14 15:49 blomadam Back to Carbon and Screens
  • 404559 12/05/14 14:42 blomadam Target Movements
  • 404555 12/05/14 12:14 blomadam Fan speed increased
  • 404554 12/05/14 11:48 blomadam Target position GUI mismatch
  • 404548 12/05/14 08:10 blomadam Target screen shot
  • Thursday

  • 404535 12/04/14 23:34 Annand Optics Run
  • 404526 12/04/14 20:16 A. Camsonne Hall A HARP scans
  • 404480 12/04/14 15:03 blomadam LHRS Q1 followup
  • 404479 12/04/14 14:16 blomadam LHRS Q1 crashed again
  • 404475 12/04/14 13:40 pking Changing the DVCS calorimeter validation parameters
  • 404474 12/04/14 13:34 pking DVCS cluster analysis from runs 10330 and 10332
  • 404473 12/04/14 12:36 hornt Q1 back on
  • 404471 12/04/14 11:16 pking LHRS Q1 crashed?
  • 404468 12/04/14 10:22 blomadam JT valve reading fluctuations
  • 404467 12/04/14 10:20 blomadam Low target coolant supply temperature
  • 404462 12/04/14 09:43 blomadam Target screen shot
  • 404431 12/04/14 04:58 akabir Target Moved
  • 404429 12/04/14 04:13 Carlos spot ++ at -1,-1 (target center?)
  • 404425 12/04/14 03:53 Carlos Unstable beam positions
  • 404424 12/04/14 03:40 adaq Low coolant temerature
  • 404423 12/04/14 01:57 Carlos Anode currents HV off
  • 404422 12/04/14 01:54 Carlos Calorimeter anode currents
  • 404415 12/04/14 00:51 Carlos BPMB x=0,y=-2.5 raster 4.5x4.5
  • 404407 12/04/14 00:03 Dien summary: swing shift 2014 Dec 03
  • Wednesday

  • 404400 12/03/14 23:41 Carlos Run 10306 spot ++
  • 404399 12/03/14 23:40 Carlos Run 10305 spot++
  • 404394 12/03/14 23:31 Dien target moved to carbon hole
  • 404393 12/03/14 22:44 Dien target moved to BeO
  • 404388 12/03/14 16:02 vmgray Day Shift: Target Summery
  • 404387 12/03/14 16:00 vmgray Target Loop 2 coolant temp
  • 404386 12/03/14 15:06 pking More about the DVCS deadtime calculation in datamon
  • 404383 12/03/14 14:42 vmgray Cryo Status Update: How does it effect you?
  • 404378 12/03/14 13:46 vmgray Target Alarm Handler Alarm Setting Changed
  • 404376 12/03/14 10:28 vmgray Raised haL2_TI_Cryo_In_R minor alarm temp
  • 404375 12/03/14 10:07 vmgray Target Alarm still going strong, temperature increasing
  • 404372 12/03/14 09:16 vmgray Target
  • 404369 12/03/14 07:34 akabir Blank GUI
  • 404366 12/03/14 03:25 akabir target temperature
  • 404365 12/03/14 01:45 akabir Target alarm
  • 404364 12/03/14 00:49 marcoapc ROC27 error when PS8>0
  • Tuesday

  • 404356 12/02/14 23:46 N. Israel target temperature chart
  • 404355 12/02/14 22:45 N. Israel target charts
  • 404349 12/02/14 18:02 jpchen target: loop 2
  • 404348 12/02/14 17:51 jpchen Target cooldown completed.
  • 404342 12/02/14 15:40 Hall A (Tech Checklist
  • 404341 12/02/14 15:40 Hall A (Tech Checklist
  • Monday